What is the domain and range for y=((x^3-4^3)/(x-4))?
Answer by stanbon(75887)

What is the domain and range for y=((x^3-4^3)/(x-4))?
------
Factored form:
---
y = [(x-4)(x^2+4x+16)]/(x-4)
-----------
There is a hole at x = 4.
So, domain is "All Real Numbers except x = 4".
------
Minimum occurs where x = -b/2a = -4/(2) = -2
f(-2) = (-2)^2+4(-2)+16) = 4-8+16 = 8
----
Range is All Real Numbers greater than or equal to 8.
=====================
Cheers,
Stan H.
===============